# Quadratic Gating Functions in Mixture of Experts: A Statistical Insight

### v2

@@ 1 Introduction
Notations. We let $[n]:=\{1,2,\ldots,n\}$ for any $n\in\mathbb{N}$. Next, for any set $S$, we denote $|S|$ as its cardinality. For any vector ${\bm{v}}\in\mathbb{R}^{d}$ and $\bm{\alpha}:=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})\in\mathbb{N}^{d}$, we let ${\bm{v}}^{\bm{\alpha}}=v_{1}^{\alpha_{1}}v_{2}^{\alpha_{2}}\ldots v_{d}^{\alpha_{d}}$, $|{\bm{v}}|:=v_{1}+v_{2}+\ldots+v_{d}$ and $\bm{\alpha}!:=\alpha_{1}!\alpha_{2}!\ldots\alpha_{d}!$, while $\|{\bm{v}}\|$ stands for its $\ell_{2}$-norm value. The probability simplex is denoted by $\Delta^{N-1}=\{{\bm{x}}\in\mathbb{R}^{N}:\sum_{i=1}^{N}x_{i}=1,x_{i}\geq 0\}$. Lastly, for any two positive sequences $(a_{n})_{n\geq 1}$ and $(b_{n})_{n\geq 1}$, we write $a_{n}=\mathcal{O}(b_{n})$ or $a_{n}\lesssim b_{n}$ if $a_{n}\leq Cb_{n}$ for all $n\in\mathbb{N}$, where $C>0$ is some universal constant. The notation $a_{n}=\mathcal{O}_{P}(b_{n})$ indicates that $a_{n}/b_{n}$ is stochastically bounded, while $a_{n}=\widetilde{\mathcal{O}}_{P}(b_{n})$ means that the previous inequality occurs up to a logarithmic factor of $n$.
2 Background
In this section, we initially introduce the MoE layer and the attention mechanism in a formal manner. Following this, we investigate the connections between the MoE framework and the attention mechanism. Furthermore, leveraging these connections, we demonstrate how quadratic gating in MoE can be understood as a self-attention mechanism.
2.1 Mixture of Experts (MoE) Layer
The mixture of experts (MoE) layer includes $N$ expert networks and a gating network. The gating network ${\bm{g}}(\cdot;\bm{\Theta}_{g})$ assigns input points to probability vectors, effectively partitioning the input space. It calculates a scoring function ${\bm{s}}(\cdot;\bm{\Theta}_{g})$ for each expert and normalizes these scores using $\bm{\sigma}(\cdot)$ to form gating probabilities ${\bm{g}}({\bm{x}};\bm{\Theta}_{g})=\bm{\sigma}({\bm{s}}({\bm{x}};\bm{\Theta}_{g}))$. Originally, dense MoE uses softmax for normalization [17], while sparse MoE uses Top-$K$ softmax for sparsity [38]. Each score $s_{i}({\bm{x}};\bm{\Theta}_{g})$ links an input ${\bm{x}}$ to an expert, and $\bm{\sigma}$ ensures that these scores sum to one. In this work, we assume that the scoring network can be written as $s_{i}({\bm{x}};\bm{\Theta}_{g})=s_{g}({\bm{x}};{\bm{\theta}}_{i})$ for all $i\in[N]$, where $s_{g}(\cdot;{\bm{\theta}})\colon\mathbb{R}^{d}\to\mathbb{R}$ is a gating scoring function. Note that this assumption holds for all commonly used gating strategies. Here, we provide a formal definition of MoE layer:
Definition 2.1 (Mixture of Experts).
Consider $N$ parameterized experts ${\bm{h}}_{i}(\cdot;{\bm{\eta}}_{i})\colon\mathbb{R}^{d}\to\mathbb{R}^{d_{o}}$ for $1\leq i\leq N$, a parameterized scoring function $s_{g}(\cdot;{\bm{\theta}})\colon\mathbb{R}^{d}\to\mathbb{R}$, and a normalization function $\bm{\sigma}\colon\mathbb{R}^{N}\to\Delta^{N-1}$. Let $\bm{\Theta}_{g}=\left({\bm{\theta}}_{1}^{\top},\ldots,{\bm{\theta}}_{N}^{\top}\right)$ and $\bm{\Theta}_{e}=\left({\bm{\eta}}_{1}^{\top},\ldots,{\bm{\eta}}_{N}^{\top}\right)$ be the set of learnable parameters for the gating network and the experts, respectively. The MoE layer is defined by
$\texttt{MoE}({\bm{x}};\bm{\Theta}_{g},\bm{\Theta}_{e}):=\sum_{i=1}^{N}g_{i}({\bm{x}};\bm{\Theta}_{g})\cdot{\bm{h}}_{i}({\bm{x}};{\bm{\eta}}_{i}),$ (4)
where gating function is given by ${\bm{g}}({\bm{x}};\bm{\Theta}_{g}):=\bm{\sigma}\left(s_{g}({\bm{x}};{\bm{\theta}}_{1}),\ldots,s_{g}({\bm{x}};{\bm{\theta}}_{N})\right)$.
2.2 Attention Mechanism
The attention mechanism [2, 41] enables transformers to focus dynamically on various parts of an input sequence, capturing essential dependencies and context. The formal definition of the attention mechanism is presented below:
Definition 2.2 (Attention Mechanism).
Consider a key matrix ${\bm{K}}=\left({\bm{k}}_{1}^{\top},\ldots,{\bm{k}}_{N}^{\top}\right)\in\mathbb{R}^{N\times d}$ that contains $N$ key vectors and a value matrix ${\bm{V}}=\left({\bm{v}}_{1}^{\top},\ldots,{\bm{v}}_{N}^{\top}\right)\in\mathbb{R}^{N\times{d_{\rm v}}}$ that includes the corresponding $N$ value vectors. We also define $s_{\alpha}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ as an attention scoring function and $\bm{\sigma}:\mathbb{R}^{N}\to\Delta^{N-1}$ as a normalization function. Given a query vector, ${\bm{q}}\in\mathbb{R}^{d}$, Attention on $({\bm{K}},{\bm{V}})$ is defined as
$\texttt{Att}({\bm{q}},{\bm{K}},{\bm{V}}):=\sum_{i=1}^{N}\alpha_{i}({\bm{q}},{\bm{K}})\cdot{\bm{v}}_{i},$ (5)
where attention weights are defined as ${\bm{\alpha}}({\bm{q}},{\bm{K}}):=\bm{\sigma}(s_{\alpha}({\bm{q}},{\bm{k}}_{1}),\ldots,s_{\alpha}({\bm{q}},{\bm{k}}_{N}))$.
The attention mechanism often uses the scaled dot product $s_{\alpha}({\bm{q}},{\bm{k}})={\langle{\bm{q}},{\bm{k}}\rangle}/{\sqrt{d}}$ for the scoring function and normalizes using softmax, simplifying to $\texttt{Att}({\bm{q}},{\bm{K}},{\bm{V}})=\mathrm{Softmax}\left(\frac{{\bm{q}}^{\top}{\bm{K}}^{\top}}{\sqrt{d}}\right){\bm{V}}$.
2.3 Mixture of Experts as an Attention
MoE models integrate attention by dynamically routing inputs to various experts via a learned gating function. The recent MoEAtt model [3] uses an attention-based routing gate to further explore the link between attention mechanisms and MoE frameworks.
The MoEAtt model functions as an attention mechanism with the query vector as the input ${\bm{x}}$, the $i$th key vector as the hidden representation of the $i$th expert for ${\bm{x}}$, and the value vectors as the outputs of the expert networks. Building on the similarity in the formulation of MoE and attention mechanisms, we now extend the MoEAtt framework to incorporate a more generalized attention-based gating mechanism:
Definition 2.3 (Attention Gating Mixture of Experts).
Extending Definition 2.1, consider parameterized query and key functions denoted by ${\bm{q}}(\cdot;\bm{\Theta}^{q})\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ and ${\bm{k}}(\cdot;\bm{\Theta}^{k}_{i})\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$, respectively, and let $s_{\alpha}$ be an attention scoring function. Then, Attention Gating Mixture of Experts is achieved by employing the following gating scoring function:
$s_{g}\left({\bm{x}};\bm{\Theta}^{q},\bm{\Theta}^{k}_{i}\right)=s_{\alpha}\left({\bm{q}}\left({\bm{x}};\bm{\Theta}^{q}\right),{\bm{k}}\left({\bm{x}};\bm{\Theta}^{k}_{i}\right)\right),$ (6)
where $\bm{\Theta}_{g}=\left\{(\bm{\Theta}^{q},\bm{\Theta}^{k}_{i}),1\leq i\leq N\right\}$ is the set of learnable gating parameters. Specifically, the output of the Attention Gating MoE can be written as
[…]
where ${\bm{k}}(\cdot;\bm{\Theta}^{k}_{i})$ represents the $i$th row of the key matrix ${\bm{K}}({\bm{x}};\bm{\Theta}^{k})$, and ${\bm{h}}_{i}({\bm{x}};{\bm{\eta}}_{i})$ represents the $i$th row of the value matrix ${\bm{V}}({\bm{x}};\bm{\Theta}_{e})$ with $\bm{\Theta}_{e}$ be the set of experts parameters.
In particular, the self-attention mechanism uses linear query and key functions along with a dot-product scoring function. This setting in Definition 2.3 leads to the formulation of a quadratic polynomial gating (2):
$g_{i}({\bm{x}};\bm{\Theta}_{g})=\frac{\exp{\left((\bm{W}^{q}{\bm{x}}+{\bm{b}}^{q})^{\top}(\bm{W}^{k}_{i}{\bm{x}}+{\bm{b}}^{k}_{i})\right)}}{\sum_{j=1}^{N}\exp{\left((\bm{W}^{q}{\bm{x}}+{\bm{b}}^{q})^{\top}(\bm{W}^{k}_{j}{\bm{x}}+{\bm{b}}^{k}_{j})\right)}},$ (8)
where $\bm{\Theta}_{g}=\{(\bm{W}^{q},{\bm{b}}^{q},\bm{W}^{k}_{i},{\bm{b}}^{k}_{i}),1\leq i\leq N\}$ denotes the collection of gating parameters. It is noteworthy that in the implementation of the self-attention mechanism, often the bias terms are omitted. Here, we can adopt this in our formulation to obtain a quadratic monomial gating MoE (3):
$g_{i}({\bm{x}};\bm{\Theta}_{g})=\frac{\exp{\left({\bm{x}}^{\top}{\bm{W}^{q}}^{\top}\bm{W}^{k}_{i}{\bm{x}}\right)}}{\sum_{j=1}^{N}\exp{\left({\bm{x}}^{\top}{\bm{W}^{q}}^{\top}\bm{W}^{k}_{j}{\bm{x}}\right)}}.$ (9)
It should be noted that commonly used gating strategies can also be seamlessly expressed as an attention gating MoE. For instance, using an identity map as the query function and a constant key function with respect to the input ${\bm{x}}$ restores the formulation to the traditional linear gating MoE.
@@ 2.3 Mixture of Experts as an Attention
Parameter count overhead. Introducing quadratic gating significantly increases the number of model parameters due to the additional quadratic terms in the gating network, which can lead to higher computational and memory demands. To mitigate this overhead, we can employ low-rank embeddings for the quadratic terms. Specifically, setting the query and key matrices $\bm{W}^{q}$ and $\bm{W}^{k}_{i}$ to dimensions $r\times d$ with $r\ll d$, we substantially reduce the number of additional parameters. This approach retains the advantages of quadratic gating while minimizing the overhead, making it a practical enhancement for MoE models. Appendix G offers a more thorough discussion on this topic.
3 Theoretical analysis of quadratic gating MoE
Motivated by the connection of quadratic MoE to attention in Section 2.3, we study the impacts of two variants of the quadratic gating on the convergence behavior of least squares expert estimation under a regression framework with the regression function taking the form of an MoE model. In particular, in Section 3.1, we examine a quadratic polynomial gating (2) in which the scoring function is a second-degree polynomial of the model input. Then, we investigate a quadratic monomial gating (3) where the scoring function is a second-degree monomial of the input in Appendix B. Furthermore, for the analysis of each quadratic gate, we derive an associated strong identifiability condition to determine which types of experts achieve better performance than others.
@@ 3.1 Quadratic Polynomial Gate
To begin with, let us formally present the regression framework used for our analysis of the quadratic polynomial gate. Assume that the data $({\bm{X}}_{1},Y_{1}),({\bm{X}}_{2},Y_{2}),\ldots,({\bm{X}}_{n},Y_{n})\in\mathbb{R}^{d}\times\mathbb{R}$ are i.i.d. sampled from the following model:
@@
where $\varepsilon_{1},\ldots,\varepsilon_{n}$ are independent Gaussian noise variables such that $\mathbb{E}[{\varepsilon_{i}}|{\bm{X}}_{i}]=0$ and $\mathrm{Var}(\varepsilon_{i}|{\bm{X}}_{i})=\sigma^{2}$ for all $1\leq i\leq n$. Additionally, we assume that ${\bm{X}}_{1},{\bm{X}}_{2},\ldots,{\bm{X}}_{n}$ are i.i.d. samples from some probability distribution $\mu$. Above, the regression function $f_{G_{*}}(\cdot)$ admits the form of a quadratic polynomial gating MoE model with $N^{*}$ experts, namely
$f_{G_{*}}({\bm{x}}):=\sum_{i=1}^{N^{*}}\frac{\exp({\bm{x}}^{\top}{\bm{A}}^{*}_{i}{\bm{x}}+({\bm{b}}^{*}_{i})^{\top}{\bm{x}}+c^{*}_{i})}{\sum_{j=1}^{N^{*}}\exp({\bm{x}}^{\top}{\bm{A}}^{*}_{j}{\bm{x}}+({\bm{b}}^{*}_{j})^{\top}{\bm{x}}+c^{*}_{j})}\cdot h({\bm{x}},{\bm{\eta}}^{*}_{i}),$ (11)
@@
where $({\bm{A}}^{*}_{i},{\bm{b}}^{*}_{i},c^{*}_{i},{\bm{\eta}}^{*}_{i})_{i=1}^{N^{*}}$ are unknown ground-truth parameters in $\mathbb{R}^{d\times d}\times\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{q}$, and $G_{*}:=\sum_{i=1}^{N_{*}}\exp(c^{*}_{i})\delta_{({\bm{A}}_{i}^{*},{\bm{b}}^{*}_{i},{\bm{\eta}}^{*}_{i})}$ denotes the associated mixing measure, a weighted sum of Dirac measures $\delta$. Meanwhile, the function $h({\bm{x}};{\bm{\eta}})$ is referred to as the expert function, which we assumed to be of parametric form.
@@
Least square estimation: To estimate the unknown parameters $({\bm{A}}^{*}_{i},{\bm{b}}^{*}_{i},c^{*}_{i},{\bm{\eta}}^{*}_{i})_{i=1}^{N^{*}}$ or, equivalently, the ground-truth mixing measure $G^{*}$, we deploy the least squares estimator [40]:
$\widehat{G}_{n}:=\operatorname*{arg\,min}_{G\in\mathcal{G}_{N}(\Theta)}\sum_{i=1}^{n}\Big{(}y_{i}-f_{G}({\bm{x}}_{i})\Big{)}^{2},$ (12)
@@
where $\mathcal{G}_{N}(\Theta):=\{G=\sum_{i=1}^{N^{\prime}}\exp(c_{i})\delta_{({\bm{A}}_{i},{\bm{b}}_{i},{\bm{\eta}}_{i})}:1\leq N^{\prime}\leq N,\ ({\bm{A}}_{i},{\bm{b}}_{i},c_{i},{\bm{\eta}}_{i})\in\Theta\}$ is the set of all mixing measures with at most $N$ components, where $N>N^{*}$. The goal of this paper is to explore the convergence properties of the estimator $\widehat{G}_{n}$ in a fixed-dimensional setting.
Given the above least squares estimator, we demonstrate in Theorem 3.1 that the convergence rate of regression estimation is parametric on the sample size.

### v3

@@ 1 Introduction
Notation. We let $[n]:=\{1,2,\ldots,n\}$ for any $n\in\mathbb{N}$. Next, for any set $S$, we denote $|S|$ as its cardinality. For any vector ${\bm{v}}\in\mathbb{R}^{d}$ and $\bm{\alpha}:=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})\in\mathbb{N}^{d}$, we let ${\bm{v}}^{\bm{\alpha}}=v_{1}^{\alpha_{1}}v_{2}^{\alpha_{2}}\ldots v_{d}^{\alpha_{d}}$, $|{\bm{v}}|:=v_{1}+v_{2}+\ldots+v_{d}$ and $\bm{\alpha}!:=\alpha_{1}!\alpha_{2}!\ldots\alpha_{d}!$, while $\|{\bm{v}}\|$ stands for its $\ell_{2}$-norm value. Additionally, for any two positive sequences $(a_{n})_{n\geq 1}$ and $(b_{n})_{n\geq 1}$, we write $a_{n}=\mathcal{O}(b_{n})$ or $a_{n}\lesssim b_{n}$ if $a_{n}\leq Cb_{n}$ for all $n\in\mathbb{N}$, where $C>0$ is some universal constant. For a sequence $(A_{n})_{n\geq 1}$ of positive random variables, the notation $A_{n}=\mathcal{O}_{P}(b_{n})$ indicates that $A_{n}/b_{n}$ is stochastically bounded, that is, for any $\epsilon>0$, there exists an $M>0$ such that $\mathbb{P}(A_{n}/b_{n}>M)<\epsilon$ for all sufficiently large $n$. Lastly, we write $A_{n}=\widetilde{\mathcal{O}}_{P}(b_{n})$ when $A_{n}=\mathcal{O}_{P}(b_{n}\log^{c}(b_{n}))$, for some $c>0$.
2 Preliminaries
In this section, we illustrate the link between self-attention mechanism and MoE in Section 2.1, and then provide background on quadratic gating MoE in Section 2.2.
2.1 Connection between MoE and Self-Attention
We start with establishing a rigorous relation between the MoE framework and the self-attention mechanism in Proposition 2.1.
Proposition 2.1.
Each row of the self-attention matrix in equation (2) can be written as an MoE with a quadratic affinity score function, which we refer to as a quadratic gating MoE.
Proof.
Firstly, recall that the self-attention matrix in equation (2) is defined as
$\mathrm{Attn}({\bm{X}})=\mathrm{Softmax}\Bigl{(}\frac{{\bm{X}}\bm{W}_{q}\bm{W}_{k}^{\top}{\bm{X}}^{\top}}{\sqrt{d_{k}}}\Bigr{)}{\bm{X}}\bm{W}_{v}\in\mathbb{R}^{N\times d_{v}},$ (3)
where ${\bm{X}}\in\mathbb{R}^{N\times d}$, $\bm{W}_{q}\in\mathbb{R}^{d\times d_{k}}$, $\bm{W}_{k}\in\mathbb{R}^{d\times d_{k}}$, and $\bm{W}_{v}\in\mathbb{R}^{d\times d_{v}}$. Let ${\bm{x}}_{i}^{\top}\in\mathbb{R}^{d}$ denote the $i$-th row of ${\bm{X}}$, and define $\bm{M}:=\frac{\bm{W}_{q}\bm{W}_{k}^{\top}}{\sqrt{d_{k}}}\in\mathbb{R}^{d\times d}$. Note that the softmax function is applied row-wise to the matrix ${\bm{X}}\bm{M}{\bm{X}}^{\top}\in\mathbb{R}^{N\times N}$. Then, the $i$-th row of the self-attention score matrix can be expressed as:
$[\mathrm{Softmax}({\bm{X}}\bm{M}{\bm{X}}^{\top})]_{i,:}=\mathrm{Softmax}({\bm{x}}_{i}^{\top}\bm{M}{\bm{X}}^{\top}).$
Thus, the $i$-th row of the self-attention output can be written as:
$[\mathrm{Attn}({\bm{X}})]_{i,:}$ $=\sum_{j=1}^{N}\mathrm{Softmax}({\bm{x}}_{i}^{\top}\bm{M}{\bm{X}}^{\top})_{j}\cdot{\bm{x}}_{j}^{\top}\bm{W}_{v}$ (4)
$=\sum_{j=1}^{N}\frac{\exp({\bm{x}}_{i}^{\top}\bm{M}{\bm{x}}_{j})}{\sum_{k=1}^{N}\exp({\bm{x}}_{i}^{\top}\bm{M}{\bm{x}}_{k})}\cdot{\bm{x}}_{j}^{\top}\bm{W}_{v}.$ (5)
Now, let $\widetilde{{\bm{x}}}=\mathrm{vec}({\bm{X}})=({\bm{x}}_{1}^{\top},\ldots,{\bm{x}}_{N}^{\top})^{\top}\in\mathbb{R}^{Nd}$ denote the vectorization of ${\bm{X}}$. For each token index $i\in{1,\ldots,N}$, let $\bm{e}_{i}\in\mathbb{R}^{N}$ denote the $i$-th standard basis vector, and define the corresponding row-selector matrix as $\bm{P}_{i}:=\bm{e}_{i}^{\top}\otimes\bm{I}_{d}$, where $\otimes$ denotes the Kronecker product, that is,
$\bm{P}_{i}=\bm{e}_{i}^{\top}\otimes\bm{I}_{d}=\begin{bmatrix}\bm{0}_{d\times d}&\cdots&\bm{0}_{d\times d}&\underbrace{\bm{I}_{d}}_{i\text{-th block}}&\bm{0}_{d\times d}&\cdots&\bm{0}_{d\times d}\end{bmatrix}\in\mathbb{R}^{d\times Nd},$
In particular, $\bm{P}_{i}$ extracts the $i$th row of ${\bm{X}}$, that is, $\bm{P}_{i}\,\widetilde{{\bm{x}}}={\bm{x}}_{i}$. Then, the $(i,j)$ entry of the logit matrix ${\bm{X}}\bm{M}{\bm{X}}^{\top}$ can be written in a quadratic form
$[{\bm{X}}\bm{M}{\bm{X}}^{\top}]_{i,j}={\bm{x}}_{i}^{\top}\bm{M}{\bm{x}}_{j}=\widetilde{{\bm{x}}}^{\top}\underbrace{\bm{P}_{i}^{\top}\bm{M}\bm{P}_{j}}_{=:\bm{B}_{ij}}\widetilde{{\bm{x}}}.$ (6)
Therefore, the $i$-th row of the self-attention matrix in equation (4) becomes:
$[\mathrm{Attn}({\bm{X}})]_{i,:}$ $=\sum_{j=1}^{N}\frac{\exp(\widetilde{{\bm{x}}}^{\top}\bm{B}_{ij}\widetilde{{\bm{x}}})}{\sum_{k=1}^{N}\exp(\widetilde{{\bm{x}}}^{\top}\bm{B}_{ik}\widetilde{{\bm{x}}})}\cdot\widetilde{{\bm{x}}}^{\top}\bm{P}_{j}^{\top}\bm{W}_{v}$
$=\sum_{j=1}^{N}\frac{\exp(\widetilde{{\bm{x}}}^{\top}\bm{B}_{ij}\widetilde{{\bm{x}}})}{\sum_{k=1}^{N}\exp(\widetilde{{\bm{x}}}^{\top}\bm{B}_{ik}\widetilde{{\bm{x}}})}\cdot\widetilde{{\bm{x}}}^{\top}\bm{E}_{j},$
where $\bm{B}_{ij}:=\bm{P}_{i}^{\top}\bm{M}\bm{P}_{j}=\frac{\bm{P}_{i}^{\top}\bm{W}_{q}\bm{W}_{k}^{\top}\bm{P}_{j}}{\sqrt{d_{k}}}\in\mathbb{R}^{Nd\times Nd}$ and $\bm{E}_{j}:=\bm{P}_{j}^{\top}\bm{W}_{v}\in\mathbb{R}^{Nd\times d_{v}}$. Hence, each row of the self-attention matrix can be represented as an MoE with quadratic affinity scores. ∎
2.2 Quadratic Gating MoE
Inspired by connection in Proposition 2.1, we consider a class of quadratic gating MoE liao2007quadratically models as an alternative to the conventional linear gating MoE. By allowing for more flexible decision boundaries through a quadratic scoring function, these models aim to provide better adaptability in expert selection. In the sequel, we particularly focus on two following quadratic gating functions:
(1) Quadratic polynomial gating function, which is inspired by the full linear embeddings with biases of keys and queries in the attention, takes the form:
${\bm{x}}\in\mathbb{R}^{d}\mapsto\mathcal{G}({\bm{x}})_{i}=\frac{\exp\left({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+{\bm{b}}_{i}^{\top}{\bm{x}}+c_{i}\right)}{\sum_{j=1}^{N}\exp\left({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+{\bm{b}}_{j}^{\top}{\bm{x}}+c_{j}\right)},$ (7)
where $({\bm{A}}_{i},{\bm{b}}_{i},c_{i})\in\mathbb{R}^{d\times d}\times\mathbb{R}^{d}\times\mathbb{R}$, for $i\in[N]$, are learnable parameters;
(2) Quadratic monomial gating function, which is motivated by the widely used linear embeddings without biases of keys and queries in the attention, is given by
${\bm{x}}\in\mathbb{R}^{d}\mapsto\mathcal{G}({\bm{x}})_{i}=\frac{\exp\left({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+c_{i}\right)}{\sum_{j=1}^{N}\exp\left({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+c_{j}\right)}.$ (8)
Subsequently, we will present a comprehensive convergence analysis for the quadratic polynomial gating of MoE in Section 3, while that for quadratic monomial gating MoE is in Section 3.2.
@@ 2.2 Quadratic Gating MoE
Parameter count overheads. Introducing quadratic gating significantly increases the number of model parameters due to the additional quadratic terms in the gating network, which can lead to higher computational and memory demands. To mitigate this overhead, we can employ low-rank embeddings for the quadratic terms. This approach retains the advantages of quadratic gating while minimizing the overhead, making it a practical enhancement for MoE models. Appendix D offers a more thorough discussion on this topic.
3 Convergence Analysis of Quadratic Gating MoE
Motivated by the link between the self-attention mechanism and quadratic gating MoE in Section 2.1, we will study the impacts of two quadratic gating functions in equations (7) and (8) on the convergence behavior of least squares expert estimators. We will examine the quadratic polynomial gate in Section 3.1, while we defer the analysis for the quadratic monomial gate to Section 3.2. Lastly, we exhibit two important practical implications on the design of self-attention and MoE models in Section 3.3.
@@ 3.1 Quadratic Polynomial Gate
To begin with, let us formally present the regression framework used for our analysis of the quadratic polynomial gate. Assume that the data $({\bm{X}}_{1},Y_{1}),({\bm{X}}_{2},Y_{2}),\ldots,({\bm{X}}_{n},Y_{n})\in\mathcal{X}\times\mathcal{Y}\subseteq\mathbb{R}^{d}\times\mathbb{R}$ are i.i.d. sampled from the following model:
@@
where $\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{n}$ are independent Gaussian noise variables such that $\mathbb{E}[{\varepsilon_{i}}|{\bm{X}}_{i}]=0$ and $\mathrm{Var}(\varepsilon_{i}|{\bm{X}}_{i})=\sigma^{2}$, for all $i\in[n]$. Additionally, we assume that ${\bm{X}}_{1},{\bm{X}}_{2},\ldots,{\bm{X}}_{n}$ are i.i.d. samples from some known probability distribution $\mu$. Above, the regression function $f_{G_{*}}(\cdot)$ admits the form of a quadratic polynomial gating MoE model with $N^{*}$ experts, namely
$f_{G_{*}}({\bm{x}}):=\sum_{i=1}^{N^{*}}\frac{\exp({\bm{x}}^{\top}{\bm{A}}^{*}_{i}{\bm{x}}+({\bm{b}}^{*}_{i})^{\top}{\bm{x}}+c^{*}_{i})}{\sum_{j=1}^{N^{*}}\exp({\bm{x}}^{\top}{\bm{A}}^{*}_{j}{\bm{x}}+({\bm{b}}^{*}_{j})^{\top}{\bm{x}}+c^{*}_{j})}\cdot\mathcal{E}({\bm{x}},{\bm{\eta}}^{*}_{i}),$ (10)
@@
where $({\bm{A}}^{*}_{i},{\bm{b}}^{*}_{i},c^{*}_{i},{\bm{\eta}}^{*}_{i})_{i=1}^{N^{*}}$ are unknown ground-truth parameters in $\mathbb{R}^{d\times d}\times\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{q}$, and $G_{*}:=\sum_{i=1}^{N_{*}}\exp(c^{*}_{i})\delta_{({\bm{A}}_{i}^{*},{\bm{b}}^{*}_{i},{\bm{\eta}}^{*}_{i})}$ denotes the associated mixing measure, that is, a weighted sum of Dirac measures $\delta$. Meanwhile, the function ${\bm{x}}\in\mathcal{X}\mapsto\mathcal{E}({\bm{x}};{\bm{\eta}})\in\mathbb{R}$ is referred to as the expert function, which we assume to be of parametric form.
@@
Least squares estimation: We estimate the unknown parameters $({\bm{A}}^{*}_{i},{\bm{b}}^{*}_{i},c^{*}_{i},{\bm{\eta}}^{*}_{i})_{i=1}^{N^{*}}$ through estimating the ground-truth mixing measure $G_{*}$ by deploying the least squares method vandeGeer-00 as follows:
$\widehat{G}_{n}:=\operatorname*{arg\,min}_{G\in\mathcal{M}_{N}(\Theta)}\sum_{i=1}^{n}\Big{(}y_{i}-f_{G}({\bm{x}}_{i})\Big{)}^{2},$ (11)
@@
where $\mathcal{M}_{N}(\Theta):=\{G=\sum_{i=1}^{N^{\prime}}\exp(c_{i})\delta_{({\bm{A}}_{i},{\bm{b}}_{i},{\bm{\eta}}_{i})}:1\leq N^{\prime}\leq N,\ ({\bm{A}}_{i},{\bm{b}}_{i},c_{i},{\bm{\eta}}_{i})\in\Theta\}$ is the set of all mixing measures with at most $N$ components, where $N>N^{*}$. The goal of this paper is to explore the convergence properties of the estimator $\widehat{G}_{n}$ in a fixed-dimensional setting.
Universal assumptions. For the sake of theory, we make the following three mild assumptions on the input and the ground-truth parameters throughout the paper.
(A.1) The input space $\mathcal{X}$ is bounded, and the parameter space $\Theta$ is compact with fixed dimension. Furthermore, the expert function ${\bm{x}}\mapsto\mathcal{E}({\bm{x}},{\bm{\eta}})$ is bounded and Lipschitz continuous with respect to ${\bm{\eta}}$.
(A.2) The last tuple of gating parameters equals to zero, that is, ${\bm{A}}^{*}_{N^{*}}=\bm{0}_{d\times d}$, ${\bm{b}}^{*}_{N^{*}}=\bm{0}_{d}$, and $c^{*}_{N^{*}}=0$. Additionally, the set of parameters $\{({\bm{A}}^{*}_{i},{\bm{b}}^{*}_{i}):i\in[N^{*}]\}$, has at least one non-zero element.
(A.3) The expert parameters ${\bm{\eta}}^{*}_{1},{\bm{\eta}}^{*}_{2},\ldots,{\bm{\eta}}^{*}_{N^{*}}$ have distinct values.
Here, the assumption (A.1) helps guarantee the convergence of the least squares estimator $\widehat{G}_{n}$ to its ground-truth counterpart $G_{*}$. Next, the first part of assumption (A.2) allows us to keep the quadratic gating MoE identifiable despite the invariance to translation of the softmax function. Meanwhile, the second part of assumption (A.2) is to maintain the dependence of the gating on the input. The assumption (A.3) is necessary to ensure that the expert values are distinct. It should be noted that these assumptions have been used in previous works chen2022theory ; nguyen2024leastsquare . Given the above assumptions, we demonstrate in Theorem 3.1 that the convergence rate of regression estimation is parametric on the sample size.